OpenStudy (anonymous):

4(2)^3 Why is the answer 32. If I follow PEMDAS, I should do the 4(2) first right?

OpenStudy (anonymous):

PEMDAS don't consider caculating series if you calcuate by PEMDAS, PEMDAS calculate by (4*2)^3 So for this calculating, you had better buy scientific calculator avavilable parenthesis function

OpenStudy (asnaseer):

PEMDAS is explained clearly here: http://www.mathsisfun.com/operation-order-pemdas.html

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

It says to do the parens first! Ok so 4x2 is 8, 8^3 is 512! Wrong answer!

OpenStudy (asnaseer):

in your particular case, it would imply you do the braces first to get:\[4(2)^3=4\times2^3\]then the exponents to get:\[4(2)^3=4\times2^3=4\times8\]

OpenStudy (asnaseer):

what you calculated was:\[(4\times2)^3=(8)^3=8^3\]

OpenStudy (anonymous):

Ok I see the () are for the ^3 not for the 4

OpenStudy (asnaseer):

yes :)

OpenStudy (asnaseer):

generally if you something like \(x*(...)^n\) it means first calculate what the \(...\) evaluates to inside the braces, then remove the braces and apply the exponent.

OpenStudy (asnaseer):

and finally multiply by x

OpenStudy (anonymous):

Thank you!

OpenStudy (asnaseer):

yw :)